# CompSpoof: A Dataset and Joint Learning Framework for Component-Level Audio Anti-spoofing Countermeasures

### v1

@@ Abstract
Component-level audio Spoofing (Comp-Spoof) targets a new form of audio manipulation where only specific components of a signal, such as speech or environmental sound, are forged or substituted while other components remain genuine. Existing anti-spoofing datasets and methods treat an utterance or a segment as entirely bona fide or entirely spoofed, and thus cannot accurately detect component-level spoofing. To address this, we construct a new dataset, CompSpoof, covering multiple combinations of bona fide and spoofed speech and environmental sound. We further propose a separation-enhanced joint learning framework that separates audio components apart and applies anti-spoofing models to each one. Joint learning is employed, preserving information relevant for detection. Extensive experiments demonstrate that our method outperforms the baseline, highlighting the necessity of separate components and the importance of detecting spoofing for each component separately. Datasets and code are available at: https://github.com/XuepingZhang/CompSpoof.
@@ 1 Introduction
Over the past decade, the ASVspoof [5, 6, 7] and other related datasets [8, 9, 10, 11] have driven significant progress in audio anti-spoofing research. Current systems [12, 13, 14, 15] typically formulate spoof detection as binary classification between bona fide and spoofed utterances. Existing anti-spoofing methods [13, 16, 17, 18] have achieved strong results under these formulations. However, these approaches implicitly assume that an utterance is either entirely genuine or entirely spoofed. The ADD challenges [19, 20] and related datasets [19, 20, 21, 22] have recently highlighted an issue of partial spoofing, where only specific time spans within an utterance are fake. However, even this setting does not address the scenario of component-level spoofing. Existing methods are unable to evaluate the genuineness of separate audio components in a mixture, which leads to poor performance when spoofing affects only one component of the audio scene.
@@ 4.1 Experimental Setup
Evaluation: We evaluate both the separation-based and baseline models on the dev and eval sets of CompSpoof, using file-level Precision, Recall, and F1 [32] as the evaluation metrics.
@@ 4.2 Experimental results and analysis
Speech
Anti-spoofing
0.860 / 0.875 / 0.863
@@
Environment
Anti-spoofing
0.846 / 0.863 / 0.849

### v2

@@ Abstract
Component-level audio Spoofing (CompSpoof) targets a new form of audio manipulation where only specific components of a signal, such as speech or environmental sound, are forged or substituted while other components remain genuine. Existing anti-spoofing datasets and methods treat an utterance or a segment as entirely bona fide or entirely spoofed, and thus cannot accurately detect component-level spoofing. To address this, we construct a new dataset, CompSpoof, covering multiple combinations of bona fide and spoofed speech and environmental sound. We further propose a separation-enhanced joint learning framework that separates audio components apart and applies anti-spoofing models to each one. Joint learning is employed, preserving information relevant for detection. Extensive experiments demonstrate that our method outperforms the baseline, highlighting the necessity of separate components and the importance of detecting spoofing for each component separately. Datasets and code are available at: https://github.com/XuepingZhang/CompSpoof.
@@ 1 Introduction
Over the past decade, the ASVspoof [15, 22] and other related datasets [18, 16, 5, 28] have driven significant progress in audio anti-spoofing research. Current systems [10, 20, 4, 6] typically formulate spoof detection as binary classification between bona fide and spoofed utterances. Existing anti-spoofing methods [20, 24, 14, 8] have achieved strong results under these formulations. However, these approaches implicitly assume that an utterance is either entirely genuine or entirely spoofed. The ADD challenges [26, 27] and related datasets [26, 27, 25, 29] have recently highlighted an issue of partial spoofing, where only specific time spans within an utterance are fake. However, even this setting does not address the scenario of component-level spoofing. Existing methods are unable to evaluate the genuineness of separate audio components in a mixture, which leads to poor performance when spoofing affects only one component of the audio scene.
@@ 4.1 Experimental Setup
Evaluation: We evaluate both the separation-based and baseline models on the dev and eval sets of CompSpoof, using file-level Precision, Recall, and F1 as the evaluation metrics.
@@ 4.2 Experimental results and analysis
Speech Anti-spoofing 0.860 / 0.875 / 0.863
@@
Environment Anti-spoofing 0.846 / 0.863 / 0.849
@@ 5 Conclusions
6 Acknowledgment
This research is funded by DKU foundation project ”Emerging AI Technologies for Natural Language Processing”. Many thanks for the computational resource provided by the Advanced Computing East China Sub-Center.